40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Jan 24 21:02:15 2011, tara, DailyProgress, NoiseBudget, RIN coupling to frequency noise vco_tf_fit.png
    Reply  Sat Apr 23 22:15:39 2011, rana, DailyProgress, NoiseBudget, RIN coupling to frequency noise 
       Reply  Mon Apr 25 10:40:00 2011, Frank, DailyProgress, NoiseBudget, RIN coupling to frequency noise 
          Reply  Wed Apr 27 22:09:10 2011, tara, DailyProgress, NoiseBudget, RIN coupling to frequency noise 
             Reply  Sun May 1 22:39:56 2011, tara, DailyProgress, NoiseBudget, RIN coupling to frequency noise TF.pngPSD.png
                Reply  Mon May 2 10:32:58 2011, Frank, DailyProgress, NoiseBudget, RIN coupling to frequency noise 
                   Reply  Tue May 3 11:59:55 2011, tara, DailyProgress, NoiseBudget, RIN coupling to frequency noise 
Message ID: 452     Entry time: Mon Jan 24 21:02:15 2011     Reply to this: 588
Author: tara 
Type: DailyProgress 
Category: NoiseBudget 
Subject: RIN coupling to frequency noise 

I'm measuring the TF between ACAV (ref) and  feedback to VCO (response).

The TF looks ok, but I still need to verify by increasing the laser power and see if the TF goes up with the same factor.

 

We are determining the coupling coefficient from intensity to frequency change via thermo-optic effect.

Using a swept sine measurement with 4000 integration cycles, from 10Hz to 100kHz, to measure the TF 

between A) ACAV RCTRANSPD (ref) and B) feedback to VCO(response), we can see some reasonable results.

I tried 2000 integration cycles with feedback to PZT, but it seems the laser is too noisy. I might try longer integration time later.

 

To make sure that the measured TF is real, we must check that when the power to the cavity is increased, the magnitude of

the TF should increase by the same factor. 

Once we have the TF between ACAV_RCTRANSPD and VCO feedback, we can calculate how RIN would couple in to frequency noise by

Thermo optic noise (RIN) [Hz] = VCO calibration [Hz/V] x TF between VCO feedback and ACAV_RCTRANSPD [V/V] x

                                                    VCO TF [V/V]    x RIN x DC level of ACAV_RCTRANSPD

 

Note the TF of the VCO box (VCO TF) is measured and fit. One pole is at 1.25 Hz, and one zero is at 38 Hz. see fig 1.

 VCO's calibration for V to Hz can be found here, The temperature is quite stable and VCO_MON has been around -0.7 to -0.3 V, we can approximate to be 1.4 MHz/V for now.

Attachment 1: vco_tf_fit.png  25 kB  | Hide | Hide all
vco_tf_fit.png
ELOG V3.1.3-